Let V be a vector space which is generated
by a finite set of vectors {v1, v2, ..., vn.}. Prove
that any linearly independent set of vectors
in V is finite and contains not more than n
elements.
1
Expert's answer
2020-03-17T13:42:38-0400

"\begin{Bmatrix}
 v_1,v_2,...v_n \\
 
\end{Bmatrix}" is spanning set of the vector space V. Let linearly independent vectors "u_1,u_2,...u_m" in vector space V and there exist "c_{ij}\isin F" such that,

"u_1=c_{11}v_1+c_{21}v_2+....+c_{n1}v_n\\
u_2=c_{12}v_1+c_{22}v_2+....+c_{n2}v_n\\
.\\
.\\
u_m=c_{1m}v_1+c_{2m}v_2+....+c_{nm}v_n\\"

Let "a_1,a_2,....a_m \isin F" be such that,


"a_1u_1+a_2u_2+....a_mu_m=0"

Therefore, this produced a homogeneous system,


"\begin{bmatrix}
 c_{11} & c_{21}...& c_{n1}\\
 c_{12} & c_{22}...&c_{n2}\\
.\\
 c_{1m} & c_{2m}...&c_{nm}
\end{bmatrix}
\begin{bmatrix}
 a_1\\
 a_2\\
.\\
a_m
\end{bmatrix}
=\begin{bmatrix}
 0\\
 0\\
.\\
0
\end{bmatrix}."

If m>n ,

since this "Ca=0" system is a under determined system,it has a non-trivial solution "\hat{a_1},\hat{a_2},...\hat{a_2}" such that "a_1u_1+a_2u_2+....a_mu_m=0" and "\hat{a_1},\hat{a_2},...\hat{a_2}" are not all zeros. Therefore "u_1,u_2,...u_m" are linearly dependent and initial linearly independent criterion is contradictory.

"\therefore" any linear independent set of vector in V is finite and not more than n.